OpenStudy (anonymous):

Stuck on this Algebraic Expression r^6 / 3r^9 I kept the 3 but i wasn't sure if i was supposed to eliminate the variable and exponent all together or keep the r and just add the exponents.

jimthompson5910 (jim_thompson5910):

assuming that "r^9" is in the denominator, you subtract the exponents

jimthompson5910 (jim_thompson5910):

so r^6/r^9 = r^(6-9) = r^(-3) then what you do is use the rule that x^(-k) = 1/(x^k) so that means r^(-3) is the same as 1/(r^3)

OpenStudy (anonymous):

1/(3r^3)

OpenStudy (anonymous):

@jim_thompson5910 its the numerator

OpenStudy (anonymous):